56. Fresh air is introduced into a room at such a rate that the air in the room is changed every 5 minutes. The room is 30 by 30 by 10 ft and the air in the room is at 70 °F with a percentage humidity of 60. Outside air at a temperature of 60 °F is used as air supply. The outside air is heated to the room temperature by a heat exchanger. How many BtU per hour must be supplied by the heat exchanger? All total pressures may be taken as 760 mmHg (A) 19500 (B) 15900 (C) 13950 (D) 9135
61) A porous material dries in the open air at a rate approximately proportional to its moisture content. If a sheet hang in the wind loses half its free moisture in the first hour, when will it have lost 85%? Weather conditions remaining the same. (A)4.72 h (B)3.24 h (C) 2.74 h (D) 7.24 h 62. A try drier is to be used for drying a solid material. Each tray is 3 ft wide by 4 ft long and the wet material has a depth of 1.8 in. The density of the wet material is 45 lb/ft. The wet material contains 3.5 Ib of water per pound of dry solid. How many trays are necessary to obtain 2500 Ib of a product containing 1 lb of water per pound of dry solid? (A)70 (B) 65 (C) 75 (D) 60
66 Soap drier reduces the water content of a given soap from 30 % by weight to 5% by weight. It is estimated that the cost of drying is PHP 10/500 kg of bone dry soap in the mixture passing through the drier. Soap containing 5wt% water sells at PHP 10/100 kg at the factory. How much should the factory charge for soap containing 30 wt% water? Express the answer as PHP per 100 kg of the moist soap containing 30 wt% water. (A)5.97 (B) 7.21 (C) 8.34 (D) 2.45
45 500 kg of air with a dry bulb of 15 °C, wet bulb of 10 °C are mixed with 1000 kg of air at 30 °C dry bulb and 15 °C wet bulb. What is the resulting temperature of the mixture? (A)70 °F (B) 85 °F (C) 25 °C (D) 21 °C
Answer #1

61.

In terms of differential equations the scenario can be written as,

-KW dt

where W is the moisture content in the material.

On integration it gives

Wo n( Kt

where Wo is the intial moisture content and W is the moisture content remaining in the solid

Given that it takes 1 hr to lose %0% of initial moisture content

Wo K 1 n 0.5 W

K = 0.693 hr-1

For losing 85% of initial moisture content

Wo n 0.15 W (0,693 * t

t = 2.74 hr

Hence option C is right

62.

Volume of wet material in each tray = 3*4*1.8*0.0833 = 1.8 ft3  

Mass of wet material in each tray = 45 * 1.8 = 81 lb

Given that 1 lb of dry solid contains 3.5 lb of water

Mass of wet solid = Mass of Dry solid + Mass of water

Mass of water = Mass of water in 1 lb of Dry solid * Mass of dry solid

Hence, 81 = Mdry + 3.5 * Mdry

Mdry​​​​​​​ = 18 lb

The product has 1 lb of water for 1 lb of dry solid.

Hence amount of dry solid = 1250 lb

As 1 tray can take 18 lb of dry solid, for 1250 trays we require 70 trays

Hence option A is right